# The Obstacle Problem Arising from the American Chooser Option

### v2

@@ 3.2. Solution to the obstacle problem
Let $V_{n}\in W^{1,2}_{p,\mathrm{loc}}(\Omega_{T}^{n})\cap C(\overline{\Omega^{n}_{T}})$ be the unique solution to the obstacle problem (3.1). Then it satisfies that
$\begin{cases}\partial_{\tau}V_{n}-\mathcal{L}V_{n}=f_{n}\quad\text{in }\ \Omega_{T}^{n},\\ \partial_{x}V_{n}(\tau,-n)=-e^{-n}\quad\text{and}\quad\partial_{x}V_{n}(\tau,-n)=e^{n}\quad\text{for }\ \tau\in[0,T],\\ V_{n}(0,x)=\max\{C(0,x),P(0,x)\}\quad\text{for }\ x\in(-n,n),\end{cases}$
[…]
$f_{n}(\tau,x)\leqslant M\quad\text{in}\ \ \Omega^{R}_{T}:=(0,T)\times(-R,R)$
@@
for some constant $M$ depending on $R$ but independent of $n$. Then it follows from $W^{1,2}_{p}$ estimates in [19, 355p] that for each small $\varrho>0,$
$\|V_{n}\|_{W^{1,2}_{p}(\Omega^{R}_{T}\setminus B_{\varrho}(0,\bar{x}))}$ $\leqslant\mathcal{K}(\|V_{n}\|_{L^{{}^{\infty}}(\Omega^{2R}_{T})}+\|C(0,\cdot)\|_{W^{2}_{p}([-2R,2R])}$
$\quad+\|P(0,\cdot)\|_{W^{2}_{p}([-2R,2R])}+\|f_{n}\|_{L^{{}^{\infty}}(\Omega^{2R}_{T})})$
$\leqslant\mathcal{K}^{\prime}$
@@
for some constant $\mathcal{K}^{\prime}$ independent of $n$. Letting $n\rightarrow\infty,$ we deduce that up to a subsequence
$V_{n}\rightharpoonup V^{R}$ $\text{in}\ \ W^{1,2}_{p}(\Omega^{R}_{T}\setminus B_{\varrho}(0,\bar{x}))\quad\text{for\ \ all }\ 1<p<\infty\ \ \text{and}\ \ \varrho>0.$
Moreover, by a standard diagonal argument, we can find a subsequence $\{V_{n_{k}}\}_{k=1}^{\infty}$ such that, as $k\rightarrow\infty$,
$V_{n_{k}}\rightarrow V^{m}$ $\text{in}\ \ C(\overline{\Omega^{m}_{T}}),$
$V_{n_{k}}\rightharpoonup V^{m}$ $\text{in}\ \ W^{1,2}_{p}(\Omega^{m}_{T}\setminus B_{\varrho}(0,\bar{x}))\quad\text{for\ \ all }\ 1<p<\infty\ \ \text{and}\ \ \varrho>0$
@@
for all large $m\in\mathbb{N}.$ Thus, $V^{m+1}=V^{m}$ in $\Omega^{m}_{T}$ so we can define $V:=V^{m}$ in $\overline{\Omega_{T}^{m}}$ for all large $m\in\mathbb{N}$. In addition, the Sobolev embedding theorem [19, Chapter II. Lemma 3.3] yields
$\partial_{x}V_{n_{k}}\rightarrow\partial_{x}V^{m}$ $\text{in}\ \ C(\overline{\Omega^{m}_{T}\setminus B_{\varrho}(0,\bar{x})}).$
Now let us prove that $V^{m}$ satisfies the following obstacle problem:
@@
$\begin{cases}\partial_{\tau}V^{m}-\mathcal{L}V^{m}\geqslant 0\quad\ \text{for }\ (\tau,x)\in\Omega^{m}_{T}\ \ \text{with}\ \ V^{m}(\tau,x)=J(\tau,x),\\ \partial_{\tau}V^{m}-\mathcal{L}V^{m}=0\quad\ \text{for }\ (\tau,x)\in\Omega^{m}_{T}\ \ \text{with}\ \ V^{m}(\tau,x)>J(\tau,x),\\ V(0,x)=J(0,x)\quad\text{for }\ x\in(-m,m).\end{cases}$
@@
Since $\partial_{\tau}V_{n}-\mathcal{L}V_{n}\geqslant 0$ in $\Omega_{T}^{n}$, by the weak convergence we see that $\partial_{\tau}V^{m}-\mathcal{L}V^{m}\geqslant 0$ and $V^{m}\geqslant J$ in $\Omega^{m}_{T}$. Thus, it remains to show that $\partial_{\tau}V^{m}-\mathcal{L}V^{m}=0$ in the set $\{V^{m}>J\}$. Fix any $\psi\in C_{c}^{\infty}(\{V^{m}>J\}\cap\Omega^{m}_{T})$ and $\varrho>0$ such that $\text{supp}(\psi)\subset\{V^{m}>J\}\cap\Omega_{T}^{m}\setminus B_{\varrho}(0,\bar{x})$. Then, there exists $\delta>0$ such that
@@
Let us fix $n\in\mathbb{N}$ and define $S_{1,{n}}$ and $S_{2,{n}}$ by
$S_{1,{n}}:=\{(\tau,x)\in\Omega^{m}_{T}:|(V_{n}-J)-(V^{m}-J)|<\delta/2\},$
$S_{2,{n}}:=\{(\tau,x)\in\Omega^{m}_{T}:|(V_{n}-J)-(V^{m}-J)|\geqslant\delta/2\}.$
Notice that $V_{n}-J>\frac{\delta}{2}$ in $S_{1,{n}}$, which implies $S_{1,{n}}\subset\{(\tau,x)\in\Omega_{T}^{n}:V_{n}(\tau,x)>J(\tau,x)\}\cap\Omega^{m}_{T}.$ Next, recall that $\partial_{\tau}V_{n}-\mathcal{L}V_{n}=0$ holds almost everywhere in $\{(\tau,x)\in\Omega_{T}^{n}:V_{n}(\tau,x)>J(\tau,x)\}$. This leads to the conclusion that
$\iint_{S_{1,n}}({\partial_{\tau}V_{n}-\mathcal{L}V_{n}})\psi=0.$
Note also that the uniform convergence of $V_{n}-J$ to $V^{m}-J$ implies that $V_{n}-J$ converges to $V^{m}-J$ in measure as $n\rightarrow\infty$. In other words, the measure of the set $S_{2,n}$ approaches 00 as $n\rightarrow\infty$. Together with $\|V_{n_{k}}\|_{W^{1,2}_{p}(\Omega^{m}_{T}\setminus B_{\varrho}(0,\bar{x}))}\leqslant M$, we deduce that
$\displaystyle{\lim_{k\to\infty}}\iint_{S_{2,n_{k}}}({\partial_{\tau}V_{n_{k}}-\mathcal{L}V_{n_{k}}})\psi=0.$
Combining the above two identities, we obtain that
$\iint_{\Omega^{m}_{T}}({\partial_{\tau}V^{m}-\mathcal{L}V^{m}})\psi$ $=\displaystyle{\lim_{k\to\infty}}\iint_{S_{1,n_{k}}\cup S_{2,n_{k}}}({\partial_{\tau}V_{n_{k}}-\mathcal{L}V_{n_{k}}})\psi=0,$
and hence $\partial_{\tau}V^{m}-\mathcal{L}V^{m}=0$ in $\{V^{m}>J\}$.
Therefore, since $V=V^{m}$ for all $m\in\mathbb{N}$, $V$ is a solution to (2.7). For the uniqueness of a solution $V$ to (2.7), we notice that by [12, Remark 3], since $|e^{x}-K|<\max\{K,1\}e^{|x|}$ for $K>0$ and $x\in\mathbb{R}$, $|\max\{C,P\}|\leqslant M_{1}e^{|x|}$ for some $M_{1}>0$ and hence $|V|\leqslant M_{2}e^{|x|}$ for some $M_{2}>0$. Therefore, by [12, Theorem 3] we obtain the uniqueness of $V$. ∎
@@ 3.3. Estimate for the derivatives of V
Let $V$ be a solution to (2.7). Then the following properties hold:
  • (1)
    $\partial_{\tau}V\geqslant 0$,
  • (2)
    $-e^{x}\leqslant\partial_{x}V\leqslant e^{x}$.
@@
(1) Fix any $\delta\in(0,T)$ and consider $V_{n,\varepsilon}(\tau+\delta,x)$ for all $(\tau,x)\in\Omega^{n}_{T-\delta}.$ Observe that $V_{n,\varepsilon}(\tau+\delta,x)$ satisfies
@@
$\begin{cases}\partial_{\tau}\{V_{n,\varepsilon}(\tau+\delta,x)-V_{n,\varepsilon}(\tau,x)\}-\mathcal{L}\{V_{n,\varepsilon}(\tau+\delta,x)-V_{n,\varepsilon}(\tau,x)\}\\ \ \ +\beta^{\prime}_{\varepsilon}(\gamma_{\tau,x})\{V_{n,\varepsilon}(\tau+\delta,x)-V_{n,\varepsilon}(\tau,x)\}=\beta^{\prime}_{\varepsilon}(\gamma_{\tau,x})\{J_{\varepsilon}(\tau+\delta,x)-J_{\varepsilon}(\tau,x)\}\ \ \text{for }(\tau,x)\in\Omega^{n}_{T-\delta},\\ \partial_{x}\{V_{n,\varepsilon}(\tau+\delta,-n)-V_{n,\varepsilon}(\tau,-n)\}=0\quad\text{and}\\ \qquad\qquad\qquad\partial_{x}\{V_{n,\varepsilon}(\tau+\delta,n)-V_{n,\varepsilon}(\tau,x)\}=0\quad\text{for }\ \tau\in[0,T-\delta],\\ V_{n,\varepsilon}(\delta,x)-V_{n,\varepsilon}(0,x)=V_{n,\varepsilon}(\delta,x)-J_{\varepsilon}(0,x)\quad\text{for }\ x\in(-n,n),\end{cases}$
where $\gamma_{\tau,x}>0$ is a positive number in between $V_{n,\varepsilon}(\tau+\delta,x)-J_{\varepsilon}(\tau+\delta,x)\geqslant 0$ and $V_{n,\varepsilon}(\tau,x)-J_{\varepsilon}(\tau,x)\geqslant 0$, determined by the mean value theorem. In the case that the latter two values are equal, we choose $\gamma_{\tau,x}$ to be large so that $\beta^{\prime}_{\epsilon}(\gamma_{\tau,x})=0$ (see (3.2)). Then from definition (3.2) it follows that $\beta^{\prime}_{\epsilon}(\gamma_{\tau,x})$ is nonnegative and bounded for $(\tau,x)\in\Omega^{n}_{T-\delta}$. In addition, by (3.4), (2.1) and Lemma 2.1 (iii), we have $(J_{\epsilon})_{\tau}=\varphi^{\prime}_{\epsilon}(C-P)(C_{\tau}-P_{\tau})+P_{\tau}\geqslant 0$. Hence the right hand side of the first equation $\beta^{\prime}_{\varepsilon}(\gamma_{\tau,x})\{J_{\varepsilon}(\tau+\delta,x)-J_{\varepsilon}(\tau,x)\}\geqslant 0$ for all $(\tau,x)\in\Omega^{n}_{T-\delta}$. Moreover, since $V_{n,\varepsilon}(\delta,x)\geqslant J_{\varepsilon}(\delta,x)$ and $V_{n,\varepsilon}(0,x)=J_{\varepsilon}(0,x)$ for all $x\in(-n,n)$, we deduce
$V_{n,\epsilon}(\delta,x)-V_{n,\epsilon}(0,x)\geqslant J_{\epsilon}(\delta,x)-J_{\epsilon}(0,x)\geqslant 0\quad\text{for }\ x\in(-n,n).$
@@
Take $\varepsilon\rightarrow 0$ and $n\rightarrow\infty$ to get the desired inequality.
(2) From the $W^{1,2}_{p}$ estimate and the Sobolev embedding theorem for the solution $V_{n,\varepsilon}$ of (3.3), we obtain that $V_{n,\varepsilon}\in C^{\frac{\alpha}{2},\alpha}(\overline{\Omega^{n}_{T}})$ for some $\alpha\in(0,1)$. Thus, the right hand side $\beta_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})\in C^{\frac{\alpha}{2},\alpha}(\Omega^{n}_{T})$. By the Schauder theory, we deduce that $V_{n,\varepsilon}\in C^{1+\frac{\alpha}{2},2+\alpha}(\Omega_{T}^{n})$. This increases the regularity of the right hand side and thus, applying the Schauder theory again yields $V_{n,\varepsilon}\in C^{2+\frac{\alpha}{2},4+\alpha}(\Omega_{T}^{n})$. By following such bootstrap argument, we conclude that $V_{n,\varepsilon}$ is smooth in $\Omega_{T}^{n}$. Differentiating the PDE (3.3) with respect to $x$, we observe that $\partial_{x}V_{n,\varepsilon}$ satisfies
@@
~$\tilde{\mathcal{L}}:=\mathcal{L}-\beta^{\prime}_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})=\frac{\sigma^{2}}{2}\partial_{xx}+(r-q-\frac{\sigma^{2}}{2})\partial_{x}-(r+\beta^{\prime}_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon}))$
Note that the zeroth-order coefficient of $\tilde{\mathcal{L}}$ being nonnegative in $\Omega^{n}_{T}$. Then, $\partial_{x}V_{n,\varepsilon}$ satisfies the equation $(\partial_{\tau}-\tilde{\mathcal{L}})(\partial_{x}V_{n,\varepsilon})=\beta^{\prime}_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})\partial_{x}J_{\varepsilon}$ in $\Omega^{n}_{T}$. Since $\mathcal{L}e^{x}=-qe^{x}$, with (3.6), we see that $\partial_{x}V_{n,\varepsilon}-e^{x}$ satisfies
@@
Thus, by the maximum principle, we deduce $\partial_{x}V_{n,\varepsilon}\leqslant e^{x}$ in $\Omega_{T}^{n}$. Similiarly, $\partial_{x}V_{n,\varepsilon}+e^{x}$ satisfies
@@
and by the maximum principle again, $\partial_{x}V_{n,\varepsilon}\geqslant-e^{x}$ in $\Omega^{n}_{T}$. Take $\varepsilon\rightarrow 0^{+}$ and $n\rightarrow\infty$ to get the desired inequality.

### v3

@@ 3.2. Solution to the obstacle problem
For each $n\in\mathbb{N}$, let $V_{n}$ be the solution to the obstacle problem (3.1) constructed in the previous subsection. By the $W^{1,2}_{p}$-regularity of $V_{n}$, we have
$\begin{cases}\partial_{\tau}V_{n}-\mathcal{L}V_{n}=f_{n}\quad\text{a.e. in }\ \Omega_{T}^{n},\\ \partial_{x}V_{n}(\tau,-n)=-e^{-n}\quad\text{and}\quad\partial_{x}V_{n}(\tau,n)=e^{n}\quad\text{for }\ \tau\in[0,T],\\ V_{n}(0,x)=\max\{C(0,x),P(0,x)\}\quad\text{for }\ x\in(-n,n),\end{cases}$
where
$f_{n}(\tau,x)=(qe^{x}-rK_{c})\mathbf{1}_{\{V_{n}=e^{x}-K_{c}\}\cap\{C>P\}}+(-qe^{x}+rK_{p})\mathbf{1}_{\{V_{n}=K_{p}-e^{x}\}\cap\{P>C\}}.$
@@
Let us fix some $m\in\mathbb{N}$. Then for each $n>2m$, $f_{n}$ is uniformly bounded in $\Omega_{T}^{m}$, with a bound depending only on $m$ and independent of $n$. Then it follows from the interior $W^{1,2}_{p}$-estimate for $1<p<\infty$ in [15, p.355] that for each small $\varrho>0,$
$\|V_{n}\|_{W^{1,2}_{p}(\Omega_{T}^{m}\setminus B_{\varrho}(0,\bar{x}))}$ $\leqslant\mathcal{K}\Big(\|V_{n}\|_{L^{\infty}(\Omega_{T}^{2m})}+\|C(0,\cdot)\|_{W^{2}_{p}((-2m,2m))}$
$\quad+\|P(0,\cdot)\|_{W^{2}_{p}((-2m,2m))}+\|f_{n}\|_{L^{\infty}(\Omega_{T}^{2m})}\Big)$
$\leqslant\mathcal{K}^{\prime},$
@@
for some constant $\mathcal{K}^{\prime}$ independent of $n$. Letting $n\to\infty$, we deduce that, up to a subsequence, there exists a function $V^{m}$ such that
$V_{n}\rightharpoonup V^{m}\quad\text{in }W^{1,2}_{p}(\Omega_{T}^{m}\setminus B_{\varrho}(0,\bar{x}))$
@@
for all $1<p<\infty$ and $\varrho>0$. In addition, the Sobolev embedding theorem [15, Chapter II. Lemma 3.3] yields
$V_{n_{k}}\rightarrow V^{m}\ \text{ in }\ C(\overline{\Omega^{m}_{T}\setminus B_{\varrho}(0,\bar{x})}).$
Combined with the continuity of $J(0,x)=\max\{C(0,x),P(0,x)\}$, it follows that $V^{m}\in C(\overline{\Omega_{T}^{m}})$ for all $m\in\mathbb{N}$. By a standard diagonal argument, we can find a subsequence $\{V_{n_{k}}\}_{k=1}^{\infty}$ such that, as $k\rightarrow\infty$,
(3.9) $V_{n_{k}}\rightharpoonup V^{m}\ \$$\text{in }\ W^{1,2}_{p}(\Omega^{m}_{T}\setminus B_{\varrho}(0,\bar{x}))\quad\text{for\ all }\ 1<p<\infty\ \text{ and }\ \varrho>0$
for each large $m\in\mathbb{N}.$ Thus, $V^{m+1}=V^{m}$ in $\Omega^{m}_{T}$ so we can define $V:=V^{m}$ in $\overline{\Omega_{T}^{m}}$ for all large $m\in\mathbb{N}$.
We now proceed to prove that $V$ is a solution to (2.7). To this end, let us fix $m\in\mathbb{N}$ and show that $V^{m}$ satisfies
@@
$\begin{cases}\partial_{\tau}V^{m}-\mathcal{L}V^{m}\geqslant 0&\text{for }(\tau,x)\in\Omega^{m}_{T}\text{ with }V^{m}(\tau,x)=J(\tau,x),\\ \partial_{\tau}V^{m}-\mathcal{L}V^{m}=0&\text{for }(\tau,x)\in\Omega^{m}_{T}\text{ with }V^{m}(\tau,x)>J(\tau,x),\\ V^{m}(0,x)=J(0,x)&\text{for }x\in(-m,m).\end{cases}$
Recall that for each $n\geqslant m$, $\partial_{\tau}V_{n}-\mathcal{L}V_{n}\geqslant 0$ a.e. in $\Omega_{T}^{m}$. Since $V_{n}\rightharpoonup V^{m}$ weakly in $W^{1,2}_{p}(\Omega_{T}^{m}\setminus B_{\varrho}(0,\bar{x}))$ for all $\varrho>0$, we observe that for any non-negative test function $\psi\in C_{c}^{\infty}(\Omega_{T}^{m})$,
$\iint_{\Omega_{T}^{m}}(\partial_{\tau}V^{m}-\mathcal{L}V^{m})\psi\,d\tau dx=\lim_{n\to\infty}\iint_{\Omega_{T}^{m}}(\partial_{\tau}V_{n}-\mathcal{L}V_{n})\psi\,d\tau dx\geqslant 0.$
Since $\psi\geqslant 0$ is arbitrary, it follows that $\partial_{\tau}V^{m}-\mathcal{L}V^{m}\geqslant 0$ a.e. in $\Omega_{T}^{m}$. Furthermore, the Sobolev embedding theorem implies that $V_{n}\to V^{m}$ in $C(\Omega_{T}^{m})$ as $n\to\infty$. Since $V_{n}\geqslant J$ in $\Omega_{T}^{m}$, passing to the limit yields
$V^{m}\geqslant J\quad\text{in }\Omega^{m}_{T}.$
Consequently, it remains to show that $\partial_{\tau}V^{m}-\mathcal{L}V^{m}=0$ on the set $\{V^{m}>J\}$.
@@
Fix any $\psi\in C_{c}^{\infty}(\{V^{m}>J\}\cap\Omega^{m}_{T})$ and $\varrho>0$ such that $\text{supp}(\psi)\subset\{V^{m}>J\}\cap\Omega_{T}^{m}\setminus B_{\varrho}(0,\bar{x})$. Then, there exists $\delta>0$ such that
@@
Let us fix $n\in\mathbb{N}$ and define $S_{n}^{(1)}$ and $S_{n}^{(2)}$ by
$S_{n}^{(1)}:=\{(\tau,x)\in\Omega^{m}_{T}:|(V_{n}-J)-(V^{m}-J)|<\delta/2\},$
$S_{n}^{(2)}:=\{(\tau,x)\in\Omega^{m}_{T}:|(V_{n}-J)-(V^{m}-J)|\geqslant\delta/2\}.$
Note that $V_{n}-J>\frac{\delta}{2}$ in $S_{n}^{(1)}$, and this implies
$S_{n}^{(1)}\subset\{(\tau,x)\in\Omega_{T}^{n}:V_{n}(\tau,x)>J(\tau,x)\}\cap\Omega^{m}_{T}.$
Next, recall that $\partial_{\tau}V_{n}-\mathcal{L}V_{n}=0$ holds almost everywhere in $\{(\tau,x)\in\Omega_{T}^{n}:V_{n}(\tau,x)>J(\tau,x)\}$. This leads to the conclusion that
$\iint_{S_{n}^{(1)}}({\partial_{\tau}V_{n}-\mathcal{L}V_{n}})\psi\,d\tau dx=0.$
Furthermore, the uniform convergence of $V_{n}-J$ to $V^{m}-J$ ensures that $V_{n}-J$ converges in measure as $n\to\infty$. In other words, the measure of the set $S_{n}^{(2)}$ tends to zero as $n\to\infty$. Together with the uniform estimate $\|V_{n_{k}}\|_{W^{2,1}_{p}(\Omega^{m}_{T}\setminus B_{\varrho}(0,\bar{x}))}\leqslant\mathcal{K}^{\prime}$, we deduce that
$\lim_{k\to\infty}\iint_{S_{n_{k}}^{(2)}}\left(\partial_{\tau}V_{n_{k}}-\mathcal{L}V_{n_{k}}\right)\psi\,d\tau dx=0.$
Combining the above two identities, we obtain
$\iint_{\Omega^{m}_{T}}({\partial_{\tau}V^{m}-\mathcal{L}V^{m}})\psi\,d\tau dx$ $=\displaystyle{\lim_{k\to\infty}}\iint_{S_{n_{k}}^{(1)}\cup S_{n_{k}}^{(2)}}({\partial_{\tau}V_{n_{k}}-\mathcal{L}V_{n_{k}}})\psi\,d\tau dx=0$
and thus, $\partial_{\tau}V^{m}-\mathcal{L}V^{m}=0$ a.e. in $\{V^{m}>J\}$. Since $m\in\mathbb{N}$ is arbitrary, we conclude that $V$ is a solution to (2.7).
For the uniqueness of a solution $V$ to (2.7), we notice that by [22, Remark 3], since $|e^{x}-K|<\max\{K,1\}e^{|x|}$ for $K>0$ and $x\in\mathbb{R}$, $|\max\{C,P\}|\leqslant M_{1}e^{|x|}$ for some $M_{1}>0$ and hence $|V|\leqslant M_{2}e^{|x|}$ for some $M_{2}>0$. Therefore, by [22, Theorem 3] we obtain the uniqueness of $V$. ∎
@@ 3.3. Estimate for the derivatives of V
Let $V$ be a solution to (2.7). Then
$\partial_{\tau}V\geqslant 0\quad\text{a.e.\ in }\ \Omega_{T}\quad\text{and}\quad-e^{x}\leqslant\partial_{x}V\leqslant e^{x}\quad\text{in }\ \Omega_{T}.$
@@
To prove the first inequality, fix some small $\delta>0$ and consider $V_{n,\varepsilon}(\tau+\delta,x)$ for all $(\tau,x)\in\Omega^{n}_{T-\delta}.$ Observe that $V_{n,\varepsilon}(\tau+\delta,x)$ satisfies
@@
$\begin{cases}\partial_{\tau}\{V_{n,\varepsilon}(\tau+\delta,x)-V_{n,\varepsilon}(\tau,x)\}-\mathcal{L}\{V_{n,\varepsilon}(\tau+\delta,x)-V_{n,\varepsilon}(\tau,x)\}\\ \ \ +\beta^{\prime}_{\varepsilon}(\gamma_{\tau,x})\{V_{n,\varepsilon}(\tau+\delta,x)-V_{n,\varepsilon}(\tau,x)\}=\beta^{\prime}_{\varepsilon}(\gamma_{\tau,x})\{J_{\varepsilon}(\tau+\delta,x)-J_{\varepsilon}(\tau,x)\}\ \ \text{for }(\tau,x)\in\Omega^{n}_{T-\delta},\\ \partial_{x}\{V_{n,\varepsilon}(\tau+\delta,-n)-V_{n,\varepsilon}(\tau,-n)\}=0\quad\text{and}\\ \qquad\qquad\qquad\partial_{x}\{V_{n,\varepsilon}(\tau+\delta,n)-V_{n,\varepsilon}(\tau,n)\}=0\quad\text{for }\ \tau\in[0,T-\delta],\\ V_{n,\varepsilon}(\delta,x)-V_{n,\varepsilon}(0,x)=V_{n,\varepsilon}(\delta,x)-J_{\varepsilon}(0,x)\quad\text{for }\ x\in(-n,n),\end{cases}$
where $\gamma_{\tau,x}>0$ is a positive number in between $V_{n,\varepsilon}(\tau+\delta,x)-J_{\varepsilon}(\tau+\delta,x)\geqslant 0$ and $V_{n,\varepsilon}(\tau,x)-J_{\varepsilon}(\tau,x)\geqslant 0$, determined by the mean value theorem. In the case that the latter two values are equal, we choose $\gamma_{\tau,x}$ to be large so that $\beta^{\prime}_{\varepsilon}(\gamma_{\tau,x})=0$ (see (3.2)). Then from definition (3.2) it follows that $\beta^{\prime}_{\varepsilon}(\gamma_{\tau,x})$ is nonnegative and bounded for $(\tau,x)\in\Omega^{n}_{T-\delta}$. In addition, by (3.4), (2.1) and Lemma 2.1 (iii), we have
$(J_{\varepsilon})_{\tau}=\varphi^{\prime}_{\varepsilon}(C-P)(C_{\tau}-P_{\tau})+P_{\tau}\geqslant 0.$
Hence, the right-hand side of the first equation,
$\beta^{\prime}_{\varepsilon}(\gamma_{\tau,x})\{J_{\varepsilon}(\tau+\delta,x)-J_{\varepsilon}(\tau,x)\},$
is nonnegative for all $(\tau,x)\in\Omega^{n}_{T-\delta}$. Moreover, since $V_{n,\varepsilon}(\delta,x)\geqslant J_{\varepsilon}(\delta,x)$ and $V_{n,\varepsilon}(0,x)=J_{\varepsilon}(0,x)$ for all $x\in(-n,n)$, we deduce
$V_{n,\varepsilon}(\delta,x)-V_{n,\varepsilon}(0,x)\geqslant J_{\varepsilon}(\delta,x)-J_{\varepsilon}(0,x)\geqslant 0\quad\text{for }\ x\in(-n,n).$
@@
Take $\varepsilon\rightarrow 0$ and $n\rightarrow\infty$, the almost everywhere differentiability of $V$ with respect to $\tau$ yields the first estimate.
Next, from the $W^{1,2}_{p}$-estimate and the Sobolev embedding theorem for the solution $V_{n,\varepsilon}$ of (3.3), we obtain that $V_{n,\varepsilon}\in C^{\frac{\alpha}{2},\alpha}(\overline{\Omega^{n}_{T}})$ for some $\alpha\in(0,1)$. Thus, the right hand side $\beta_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})$ belongs to $C^{\frac{\alpha}{2},\alpha}(\Omega^{n}_{T})$. By standard Schauder theory, we deduce that $V_{n,\varepsilon}\in C^{1+\frac{\alpha}{2},2+\alpha}(\Omega_{T}^{n})$. This enhances the regularity of the right-hand side, and thus, a subsequent application of Schauder theory yields $V_{n,\varepsilon}\in C^{2+\frac{\alpha}{2},4+\alpha}(\Omega_{T}^{n})$. Iterating this bootstrap argument, we conclude that $V_{n,\varepsilon}$ is smooth in $\Omega_{T}^{n}$. This smoothness justifies differentiating the equation in (3.3) with respect to $x$. Consequently, we see that $\partial_{x}V_{n,\varepsilon}$ satisfies
@@
$\tilde{\mathcal{L}}:=\mathcal{L}-\beta^{\prime}_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})=\frac{\sigma^{2}}{2}\partial_{xx}+(r-q-\frac{\sigma^{2}}{2})\partial_{x}-(r+\beta^{\prime}_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon}))$
Then, $\partial_{x}V_{n,\varepsilon}$ satisfies the equation $(\partial_{\tau}-\tilde{\mathcal{L}})(\partial_{x}V_{n,\varepsilon})=\beta^{\prime}_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})\partial_{x}J_{\varepsilon}$ in $\Omega^{n}_{T}$. Since $\mathcal{L}e^{x}=-qe^{x}$, with (3.6), we see that $\partial_{x}V_{n,\varepsilon}-e^{x}$ satisfies
@@
Note that the zeroth-order coefficient of $\tilde{\mathcal{L}}$ is nonnegative. Thus, by the maximum principle, we deduce $\partial_{x}V_{n,\varepsilon}\leqslant e^{x}$ in $\Omega_{T}^{n}$. Similarly, $\partial_{x}V_{n,\varepsilon}+e^{x}$ satisfies
@@
and again by the maximum principle, it follows that $\partial_{x}V_{n,\varepsilon}\geqslant-e^{x}$ in $\Omega^{n}_{T}$. Note each weak convergence (3.7), (3.9) with $p>3$ and the Sobolev embedding theorem [15, Chapter II. Lemma 3.3] to see that $\partial_{x}V_{n,\varepsilon}\to\partial_{x}V$ uniformly as $\varepsilon\rightarrow 0^{+}$ and $n\rightarrow\infty$. Thus, taking $\varepsilon\to 0^{+}$ and $n\to\infty$ yields the result.